Find the limit
x tends to 2
(x^2+4x–1)/(x^2–2x)
1
Expert's answer
2021-09-20T08:18:33-0400

"\lim{x\ \rightarrow2}{\frac{x^2+4x-1}{x^2-2\ x}}"

"\lim{x\ \rightarrow2}{\frac{\frac{x^2}{x^2}+\frac{4x}{x^2}-\frac{1}{x^2}}{\frac{x^2}{x^2}-\frac{2x}{x^2}}}"

"\frac{1+\frac{4}{2}-\frac{1}{4}}{1-\frac{2}{2}}=\frac{\frac{5}{4}}{0}\"

The limit is undefined